OpenStudy (anonymous):

How many number of handshakes are possible if there are 25 persons and each one will shake each others hand?

OpenStudy (anonymous):

didnt i just answer this?

OpenStudy (anonymous):

24+23+22 and so on.. can you think of the reason why?

OpenStudy (anonymous):

there will be 300 shakehands. becuase first person will shake with 24 persons. secong d person will shake with left out i.e 23 3rd 22 and so on...... so the total no. of shakehands= n(n+1)/2 =24 * 25/2 =300

OpenStudy (anonymous):

wow thank you so much...it really helps..